develop a photon language programming code for cloud and Lene Dev ECR refrigeration and air conditioning system hybrid connected with dehumidifier which produces 60 50 and 45 degree centigrade temperature also write the parameter which are affecting the performance of the system developed also claim the temperature attained by this time
Answer #1

"Input data"

T[8] = 300 [K] "Gas compressor inlet"

P[8] = 14.7 [kPa] "Assumed air inlet pressure"

"Pratio = 14" "Pressure ratio for gas compressor"

T[10] = 1400 [K] "Gas turbine inlet"

T[12] = 460 [K] "Gas exit temperature from Gas-to-steam heat exchanger "

P[12] = P[8] "Assumed air exit pressure"

W_dot_net=450 [MW]

Eta_comp = 1.0

Eta_gas_turb = 1.0

Eta_pump = 1.0

Eta_steam_turb = 1.0

P[5] = 8000 [kPa] "Steam turbine inlet"

T[5] =(400+273) "[K]" "Steam turbine inlet"

P[6] = 600 [kPa] "Extraction pressure for steam open feedwater heater"

P[7] = 20 [kPa] "Steam condenser pressure"

"GAS POWER CYCLE ANALYSIS"

"Gas Compressor anaysis"

s[8]=ENTROPY(Air,T=T[8],P=P[8])

ss9=s[8] "For the ideal case the entropies are constant across the compressor"

P[9] = Pratio*P[8]

Ts9=temperature(Air,s=ss9,P=P[9])"Ts9 is the isentropic value of T[9] at compressor exit"

Eta_comp = w_gas_comp_isen/w_gas_comp "compressor adiabatic efficiency, w_comp > w_comp_isen"

h[8] + w_gas_comp_isen =hs9"SSSF conservation of energy for the isentropic compressor, assuming: adiabatic,

ke=pe=0 per unit gas mass flow rate in kg/s"

h[8]=ENTHALPY(Air,T=T[8])

hs9=ENTHALPY(Air,T=Ts9)

h[8] + w_gas_comp = h[9]"SSSF conservation of energy for the actual compressor, assuming: adiabatic,

ke=pe=0"

T[9]=temperature(Air,h=h[9])

s[9]=ENTROPY(Air,T=T[9],P=P[9])

"Gas Cycle External heat exchanger analysis"

h[9] + q_in = h[10]"SSSF conservation of energy for the external heat exchanger, assuming W=0, ke=pe=0"

h[10]=ENTHALPY(Air,T=T[10])

P[10]=P[9] "Assume process 9-10 is SSSF constant pressure"

Q_dot_in"MW"*1000"kW/MW"=m_dot_gas*q_in

"Gas Turbine analysis"

s[10]=ENTROPY(Air,T=T[10],P=P[10])

ss11=s[10] "For the ideal case the entropies are constant across the turbine"

P[11] = P[10] /Pratio

Ts11=temperature(Air,s=ss11,P=P[11])"Ts11 is the isentropic value of T[11] at gas turbine exit"

Eta_gas_turb = w_gas_turb /w_gas_turb_isen "gas turbine adiabatic efficiency, w_gas_turb_isen > w_gas_turb"

h[10] = w_gas_turb_isen + hs11"SSSF conservation of energy for the isentropic gas turbine, assuming:

adiabatic, ke=pe=0"

hs11=ENTHALPY(Air,T=Ts11)

h[10] = w_gas_turb + h[11]"SSSF conservation of energy for the actual gas turbine, assuming: adiabatic,

ke=pe=0"

T[11]=temperature(Air,h=h[11])

s[11]=ENTROPY(Air,T=T[11],P=P[11])

"Gas-to-Steam Heat Exchanger"

"SSSF conservation of energy for the gas-to-steam heat exchanger, assuming: adiabatic,

W=0, ke=pe=0"

m_dot_gas*h[11] + m_dot_steam*h[4] = m_dot_gas*h[12] + m_dot_steam*h[5]

h[12]=ENTHALPY(Air, T=T[12])

s[12]=ENTROPY(Air,T=T[12],P=P[12])

"STEAM CYCLE ANALYSIS"

"Steam Condenser exit pump or Pump 1 analysis"

Fluid$='Steam_IAPWS'

P[1] = P[7]

P[2]=P[6]

h[1]=enthalpy(Fluid$,P=P[1],x=0) {Saturated liquid}

v1=volume(Fluid$,P=P[1],x=0)

s[1]=entropy(Fluid$,P=P[1],x=0)

T[1]=temperature(Fluid$,P=P[1],x=0)

w_pump1_s=v1*(P[2]-P[1])"SSSF isentropic pump work assuming constant specific volume"

w_pump1=w_pump1_s/Eta_pump "Definition of pump efficiency"

h[1]+w_pump1= h[2] "Steady-flow conservation of energy"

s[2]=entropy(Fluid$,P=P[2],h=h[2])

T[2]=temperature(Fluid$,P=P[2],h=h[2])

"Open Feedwater Heater analysis"

y*h[6] + (1-y)*h[2] = 1*h[3] "Steady-flow conservation of energy"

P[3]=P[6]

h[3]=enthalpy(Fluid$,P=P[3],x=0) "Condensate leaves heater as sat. liquid at P[3]"

T[3]=temperature(Fluid$,P=P[3],x=0)

s[3]=entropy(Fluid$,P=P[3],x=0)

"Boiler condensate pump or Pump 2 analysis"

P[4] = P[5]

v3=volume(Fluid$,P=P[3],x=0)

w_pump2_s=v3*(P[4]-P[3])"SSSF isentropic pump work assuming constant specific volume"

w_pump2=w_pump2_s/Eta_pump "Definition of pump efficiency"

h[3]+w_pump2= h[4] "Steady-flow conservation of energy"

s[4]=entropy(Fluid$,P=P[4],h=h[4])

T[4]=temperature(Fluid$,P=P[4],h=h[4])

w_steam_pumps = (1-y)*w_pump1+ w_pump2 "Total steam pump work input/ mass steam"

"Steam Turbine analysis"

h[5]=enthalpy(Fluid$,T=T[5],P=P[5])

s[5]=entropy(Fluid$,P=P[5],T=T[5])

ss6=s[5]

hs6=enthalpy(Fluid$,s=ss6,P=P[6])

Ts6=temperature(Fluid$,s=ss6,P=P[6])

h[6]=h[5]-Eta_steam_turb*(h[5]-hs6)"Definition of steam turbine efficiency"

T[6]=temperature(Fluid$,P=P[6],h=h[6])

s[6]=entropy(Fluid$,P=P[6],h=h[6])

ss7=s[5]

hs7=enthalpy(Fluid$,s=ss7,P=P[7])

Ts7=temperature(Fluid$,s=ss7,P=P[7])

h[7]=h[5]-Eta_steam_turb*(h[5]-hs7)"Definition of steam turbine efficiency"

T[7]=temperature(Fluid$,P=P[7],h=h[7])

s[7]=entropy(Fluid$,P=P[7],h=h[7])

"SSSF conservation of energy for the steam turbine: adiabatic, neglect ke and pe"

h[5] = w_steam_turb + y*h[6] +(1-y)*h[7]

"Steam Condenser analysis"

(1-y)*h[7]=q_out+(1-y)*h[1]"SSSF conservation of energy for the Condenser per unit mass"

Q_dot_out*Convert(MW, kW)=m_dot_steam*q_out

"Cycle Statistics"

MassRatio_gastosteam =m_dot_gas/m_dot_steam

W_dot_net*Convert(MW, kW)=m_dot_gas*(w_gas_turb-w_gas_comp)+ m_dot_steam*(w_steam_turb -

w_steam_pumps)"definition of the net cycle work"

Eta_th=W_dot_net/Q_dot_in*Convert(, %) "Cycle thermal efficiency, in percent"

Bwr=(m_dot_gas*w_gas_comp + m_dot_steam*w_steam_pumps)/(m_dot_gas*w_gas_turb +

m_dot_steam*w_steam_turb) "Back work ratio"

W_dot_net_steam = m_dot_steam*(w_steam_turb - w_steam_pumps)

W_dot_net_gas = m_dot_gas*(w_gas_turb - w_gas_comp)

NetWorkRatio_gastosteam = W_dot_net_gas/W_dot_net_steam